ASSUME ALL CORPORATIONS ARE C CORPORATIONS

Corporation X-a C Corporation-distributed the following to its sole shareholder-John-during the year:
Inventory-cost/basis of $5,000 with a fair market value of $9,000.
The following facts are provided:
Corp X has accumulated E and P of $ 12, 000 and current E and P of $2,000 - without regard to this distribution. Corp X's tax rate is 21 percent. John has a basis in the corporate stock of $23,000.
1. What are the tax effects of this distribution to the corporation?
2. What is the Accumulated earnings and profits as of the end of the year?
3.what are the effects of this distribution to John?
4.what is John's basis in his Corp. X stock after the receipt of this distribution?
Answer #1

1) A corporation will not recognize any gain or loss on a distribution of cash to its shareholders.[13] But if the corporation distributes appreciated property, the corporation must recognize gain as if the property were sold to the shareholder at fair market value

2)Accumulated earnings and profits- 12,000-2000=10,000

Tax rate-10,000*21/100=2,100

3) There is no gain or loss that will be recognized on the distribution of inventory to John. John has a zero basis in of Rs 9000 in the inventory at face value of market price and in which he has a basis of Rs 5000 cash distribution,

The distribution reduces his basis in the partnership interest to(10000-9000-5000=-4000). Since cash distribution exceeds John's basis,John will therefore recognize a gain for the excess.

4) If a redemption is to be carried out on a deferred payment basis, the inactive spouse can defer tax by using the installment method. However, if the redemption is taxed as a dividend to the active spouse, installment sale treatment is not available